How many grams of H2O are needed if 41.85g of O2 are to be produced from the reaction below? Show the solution




2K2O2 + 2H2O -----> 4KOH + O2

1
Expert's answer
2022-06-01T01:14:04-0400

Mr (H2O) = 18 g/mol

Mr (O2) = 32 g/mol


2K2O2 + 2H2O = 4KOH + O2

According to the above reaction, 2 mol (36 g) of H2O reacts to form 1 mol (32 g) of O2. We find how many g of H2O must be reacted to form 41.85 g of O2 :

36g H2O —> 32g O2

X —> 41,85g O2

X = 36 * 41,85 / 32 = 47,08 g

ANSWER: 47,08 g H2O